OpenStudy (anonymous):

The area of a rectangle can be found by multiplying the length by the width. A rectangular postage stamp has a length of nine-tenths inch and a width of five-sixths inch. What is the area of the stamp, in square inches?

OpenStudy (anonymous):

\[Area = Length \times Width\]\[Area = \frac{ 9 }{ 10 } \times \frac{ 5 }{ 6 }\]\[Area = \frac{ 45 }{ 60 }\]\[Area = 0.75 in ^{2}\]

OpenStudy (anonymous):

thanks dude! :)

OpenStudy (anonymous):

Your welcome.